You are eyeing an investment in a corporate bond which has a YTM of 15.00%, and a stated rate of interest of 14.00% with a maturity of 2 years. What is the price of this bond? (I am struggling to try and solve this on my ti-84 calculator, so if you could show how to solve it on that, that would be great.)

a. $983.25

b. $2,805.09

c. $1,217.71

d. $869.94
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =2x2
Bond Price =∑ [(14*1000/200)/(1 + 15/200)^k]     +   1000/(1 + 15/200)^2x2
                   k=1
Bond Price = 983.25
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %/coupons per year=1000*14/(2*100)
I/Y =15/2
N =2*2
FV =1000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(15/(2*100),2*2,-14*1000/(2*100),-1000,)